OpenStudy (anonymous):

Can someone please explaine to me how to do this problem step by step? Solve the equation or formula for the indicated variable. S = 5r²t, for t

OpenStudy (anonymous):

S/5r^2 = t

OpenStudy (anonymous):

But these are my answer choices....attached below

OpenStudy (anonymous):

for further explanation, you just simply divide both sides by 5r^2 to cancel 5r^2 on the right side. OR you can multiply both sides by 1/5r^2, it has the same results anyway. S = 5r^2t S/5r^2 = t or t = S/5r^2

OpenStudy (anonymous):

Thank you very much!!!:)